FEL Upgrade Project Weekly Brief June 3-7, 2002
On Wednesday the installation crew held a "grouting" party. With the assistance of many of the FEL team members who dressed for their part, all of the stands which had previously been surveyed and installed for FEL Upgrade hardware were grouted to the floor. Subsequently, the vault was cleaned up in preparation for next weeks scheduled warm-up of the injector cryounit for alignment checks and replacement of a leaking beamline valve.
The review process continues. This week we received draft comments from our seminanaual review held on May 1-3 for our review and comment. We thank our hard working review team (and the court reporter-John Albertine) for their time and effort spent on the review and report.
We prepared dry run presentations for our upcoming annual DOE review of Jefferson Labs Science and Technology scheduled for July 15-16.
WBS 4 (Injector):
The high voltage power supply in the W&M implantation has been repaired after the high voltage connector failed. A test run with a sample electrode inside the plasma proved that the system is working properly now. The shielding door for the ball cathode will be implanted this afternoon. The support tube which suffered arc damage in the last run has had the implant layer etched away and is being polished. The tapered shielded bellows drawings are in sign off, and the bellows for it have been received. The shielded bellows/BPM unit has had the BPM tested and the design is now finalized, ordered bellows for the unit.
Gun HVPS - The Gun HVPS Conditioning Resistor was fired this week. After looking at the results and then sending photo's to the vendor of the resistive material, it was determined the resistor was at the firing temperature for too long a time. It is not possible to reduce the firing time, so a lower temperature was suggested. The resistive material formed small unconnected puddles on the silver contacts. The resistor section looked like an open circuit above 1 kM. This resistor will be used to determine a better firing profile. It was physically broken during handling, but it was already unusable due to the over-firing. The next firing at a lower temperature will be done early next week.
WBS 6 (RF):
Quarter HVPS - The wiring for unit 3 is finished and unit 4 is nearly ready also. The PLC code is due by June 21st.
Quarter Klystrons - The 2nd and 3rd klystron cart was returned to the Machine Shop and are due out next week. One of the changes found on the 1st cart did not make it into the drawing package. The correction to the drawing has already been completed.
WBS 8 (Instrumentation):
Progress is made on the interface for the beam viewer chassis, assembly of Beam Loss Monitor heads, and fabrication of 8 channel stepper motor drives.
Most of the group attended the group grouting party and subsequent clean up. The ion implantation station went down due to a cable fault and a damaged HV/oil feedthrough. A number of problems were cleared up and the system is up and running again.
Production issues were worked through with assembly of quad girders. The short stripline BPM for the injector was tested after welding in of the feedthroughs. There was a problem with one of the four, it is being repaired.
WBS 9 (Transport):
Optical Chicane Dipoles (GW)
At Magnet Enterprises International (MEI) in Oakland CA. The Epoxy manufacturer claims that the residual oil-like material on the coil surface after potting is a residual from the formulation that should be there and is not harmful. They have verified epoxy strength of the sample material we sent to them. This matches our Durometer readings showing full strength to the epoxy.
We still want to track down the source of the hazing (perhaps micro de-bonding of the glass fibers of the outer ground wrap) that all coils using NMA hardener from this manufacturer have had for some time. So far we believe this is a cosmetic flaw. (Note that may coil customers opt to paint the coils or include an opaque filler to gloss over these cosmetic problems. We rather believe in WYSIWYG.)
The first batch of core plates for the 4 dipoles of the IR Machine were completed. The first article was assembled ready for gap check today.
Cores for the second batch of 6 UV Magnets remain in final machining.
Injector Dipoles (DU/DV)
At Master Machine, two sets of GV upper and lower pieces are complete and others remain in machining. The return legs will go for final grinding next week. Shipping will be in two weeks.
At Wang NMR, GU/GV coil potting starts next week and will finish in three weeks, just in time for cores from Master Machine to start Assembly. These will be the first dipoles delivered - in five to six weeks
Arc 180 Degree Dipoles (GY)
Assembled 2nd Core looks good at Bosma. I was not able to measure the gap because the measuring instrument was in my luggage which was delayed from the airplane flight. We will send our instrument next week for them to measure themselves.
(I also inspected the first core and all parts that had arrived at Wang NMR on my visit the of 6/6/02. The parts look good and not damaged.
At Wang NMR, the second GY Coil (of four) is in winding its last layer much better than first. They improved and reworked the tool substantially they have instant feedback to a skilled tooling machinest.
GY Coil Potting fixture now designed. They expect ten-day turn around on its manufacture. Coils will be ready by end of August. We have to work shim glue down in parallel.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
Process Equipment Co. (PECo) continued manufacturing the GQ-GX Cores.
Half the GX cores are through heat treat.
Commercial parts and other small parts are on order or in fabrication, many are delivered or complete.
Nothing was discussed about the excessive material to be removed for rough machining. I believe they are still putting together a claim.
Their plan for first article GQ on June 19 looks good. Their delivery will match coil availability.
They are fully responsive to Rev A that adds the shimming feature to the pole tip edges.
At Wang NMR, the second winding line is being set up for GQ-GX Coils. They will start winding late next week. They plan to deliver these magnets second.
3 inch quad (QX)
The last shipment of QX magnets (for the UV branch) was sent on Wednesday.
We have finally resolved our magnet measurement issues and production measurement of all the QX quads resumed on an 18 hr per day basis.
Trim Quad (QT)
Milhous Control of Virginia has all coils wound. The second magnet is ready for assembly.
Tom Hiatt has resolved some magnet model problems and is about to start adding the GC corrector to the trim quad to see if it a viable option.
DULY Research is nearly finished with the drawings of the magnet.
Work on this magnet is on furlough until the sextupole is designed.
Beam Line and Vacuum
The procurement documents for the ARC Chambers are nearing signoff.
The X Chamber has started fabrication in the Shop.
All pedestals are installed that are here and will not interfere with future installation
A number of girders are in the enclosure awaiting installation.
The drawings for the recirculation region in front of the Cryomodule are near sign-off.
Design of the optical chicane region with its chamber continues.
Design of the regions around the wiggler continues.
WBS 11 (Optics):
Progress on the upgrade occurred in the following areas:
We have completed fabrication of the second optical cavity assembly and test stand, and are awaiting receipt of the first parts (leadscrew fixtures, etc), due later this month. The vendor for the OCMMS beamline parts has informed us they will not deliver the parts until next week, nearly one month late. While there is sufficient schedule float to not be an issue, the same vendor is making a sizeable portion of the optical cavity internals, and we do want to know whether they will make their schedule. The OCMMS external mirrors were inspected and found to meet specs. The drift distance from the external OCCMS mirror to the center of the wiggler was determined, and will be considered in making the optical layout. We received a budgetary estimate for optical tables to support the OCMMS hardware. Given the cost, we are evaluating the use of simpler breadboard-style optical benches. The first tests using Ga:In eutectic as a heat conduction compound showed that a gap of 0.050" was too large. We are going to try again next week with a much narrower gap. So, far, tests of the mirror mounted with the improved In:Ag brazing technique continue to look very good.
We anticipate brazing of the first article 50kW optical beam dump within the next two weeks. This is a continued slip from our original date, but this item has considerable schedule float. Given the unwillingness of on vendor of metal foam to sell us their material, we are checking with other vendors. As an alternative to the black copper oxide coating we are considering, the anticipated backscattered radiation level using Molectron's flame-sprayed coating (which we know worked well for our 15 kW beam dumps) was modeled. The most probable number of reflections is ~ 4, and with the manufacturer's value for the reflectance (4%) we make our backscatter specification.
Our compact spectrometer arrived this week. It will be used as a pulsewidth monitor (for our ultrafast laser and the compressed output of the drive laser) and for spectroscopic analyses in the near IR. The ultrafast laser oscillator shipped on May 31st. We anticipate it arriving today (6/7) or early next week.
We received parts for the prototype backplane cooled mirror mount and conveyor chain that will be used for routing of the water lines in the mirror cassettes in the User Labs. Some brazing work remains on the backplane cooled mount, and then it will be checked in the Mirror Test Stand. We are also performing some temporary modifications to the UL 3 mirror cassette in order to perform particulate studies. These are needed to certify our UHV lubrication procedures.
M. Shinn attended a conference on laser advanced materials processing in Osaka, Japan, where she presented an invited talk on the FEL. It was well-received, and generated some interest from industry representatives.